OpenStudy (anonymous):

Calculate the area of triangle ABC with altitude CD, given A (-7, -1), B (-1, 5), C (0, 0), and D (-3, 3).

OpenStudy (anonymous):

9 square units 18 square units 18.5 square units 21 square units

OpenStudy (anonymous):

@mathstudent55

OpenStudy (accessdenied):

Did you figure this one out (hence closure)?

OpenStudy (anonymous):

No @AccessDenied

OpenStudy (accessdenied):

Alright. Well, you want to find Area. That entails knowing two things to apply this area formula: \( A_{\text{triangle}} = \frac{1}{2} b h \) The base, and the height.

OpenStudy (anonymous):

Yes.

OpenStudy (accessdenied):

So, we have the three vertices of the triangle. We also have this extra D coordinate which gives an altitude. According to the definition of an altitude, though, that is pretty much the height of the triangle assuming the side perpendicular is the base, correct?

OpenStudy (anonymous):

Yes. Correct.

OpenStudy (accessdenied):

So, if you have trouble visualizing which side is perpendicular, I would recommend simply plotting the points and seeing it for yourself. But CD is our altitude and also our height. The side opposite of vertex C is the segment AB.

OpenStudy (accessdenied):

...Okay that's also good. :p

OpenStudy (anonymous):

Okay.

OpenStudy (accessdenied):

So the final goal is to find the length of CD, the height, and AB, our base. Then they are plugged right in to our Area formula.

OpenStudy (anonymous):

CD ; 4.25

OpenStudy (anonymous):

AB ; 8.50

OpenStudy (accessdenied):

I am not familiar with GeoGebra or if it allows you to find the lengths freely. Otherwise, distance formula between the two points works too. In either case, though, whichever is appropriate for your needs. :)

OpenStudy (accessdenied):

So the Area is then: A = 1/2 * AB * CD

OpenStudy (anonymous):

18...

OpenStudy (anonymous):

Correct?

OpenStudy (accessdenied):

18.0625.. seems close enough to 18 to not be 18.5.

OpenStudy (accessdenied):

So, correct. (Had to check with distance formula to make sure... with the exact measurements, you get exactly 18)

OpenStudy (anonymous):

Yeah thats what i got thanks! I have some more can you help?

OpenStudy (anonymous):

Find the perimeter of a quadrilateral with vertices at R (-2, 1), S (-5, 5), T (2, 5), U (5, 1). Round your answer to the nearest hundredth when necessary. 22 units 24 units 28 units 36 units

OpenStudy (accessdenied):

Perimeter means the length around the figure. You have the coordinates, and you want to find the lengths or distances of each pair of points creating a side of the quadrilateral.

OpenStudy (anonymous):

Yes i got that.

OpenStudy (accessdenied):

You might be alright to assume it goes in order: R connects to S, S connects to T, T connects to U, and U loops back on R. A plot will confirm whether that is true or not.

OpenStudy (anonymous):

24 units

OpenStudy (accessdenied):

Looks good to me. Indeed, that assumption is true; RSTU is in order.

OpenStudy (anonymous):

For the line segment whose endpoints are A (0, 0) and B (4, 3), find the x value for the point located 2/3 the distance from A to B. 2.2 2.7 3.3 3.5

OpenStudy (accessdenied):

Well, it would be good to start with finding the distance. Then we can take 2/3 of it for the length of this segment: 

OpenStudy (anonymous):

2.7..?

OpenStudy (accessdenied):

2.7 seems to be correct. Those would make for similar triangles, so hypotenuse ratio = x-length ratio 4/x = 5/ (5 (2/3)) x = 4 * 2/3

OpenStudy (anonymous):

Point L is located at (4, -3) and M is located at (-8, 5). Find the y value of the point that lies halfway between L and M. 1.5 4 3.5 1

OpenStudy (accessdenied):

So, we want the midpoint of L and M. You know how to find the midpoint of two given points?

OpenStudy (anonymous):

Yes i just can't remember at the moment..

OpenStudy (accessdenied):

The easiest way to remember it is the average of the corresponding coordinates. Add the two x-values and divide by 2 for the midpoint x-coordinate Add the two y-values and divide by 2 for the midpoint y coordinate

OpenStudy (anonymous):

Wow thats a great way to remember actually

OpenStudy (anonymous):

Answer is 1..?

OpenStudy (accessdenied):

Correct. :)

OpenStudy (anonymous):

A segment with endpoints I (5, 2) and J (9, 10) is divided by a point K such that IK and IJ form a 2:3 ratio. Find the y value for K. 4.6 5.4 4.8 5.2

OpenStudy (accessdenied):

So, the two sides make the ratio: IK / IJ = 2 / 3 We could find the distance between K(x, y) and I(5, 2) along with the distance between K(x, y) and J(9, 10). Then we could intersect with the line through the two points I and J. This problem seems more involved...

OpenStudy (accessdenied):

In both cases, you would use distance formula of the two points: \( d = \sqrt{ (x_2 - x_1)^2 + (y_2 - y_1)^2} \)

OpenStudy (anonymous):

input them?

OpenStudy (accessdenied):

Hang on, I was thinking KI and KJ. Sorry. Let me rethink that.

OpenStudy (anonymous):

sure no problem..

OpenStudy (accessdenied):

So the statement IK : IJ = 2 : 3 implies that IK is 2/3 the length of IJ. So we are finding the point K whose distance from I is 2/3 IJ.

OpenStudy (anonymous):

Yes.

OpenStudy (accessdenied):

So we would still find the length of IJ which would lead us the length of IK. Then we just need to find the point 2/3 IJ away from I. The geometric argument is that we are swiping a circle of radius 2/3 IJ with center at I around and trying to find the place where the line IJ intersects the circle.